# Paravortices: loop braid representations with both generators involutive

### v1

@@ 4. Mixed doubles representations
(4.2) The category $\mathsf{MD}$} is defined as follows. It is the strict monoidal category of mixed doubles groups ${\mathsf{MD}}_{n}$, analogous to the braid category or the Sym category. The object monoid is $(\mathbb{N}_{0},+)$, thus it is freely generated by the object 1. The monoidal product on morphisms is given by the inclusion $\phi:{\mathsf{MD}}_{m}\times{\mathsf{MD}}_{n}\rightarrow{\mathsf{MD}}_{m+n}$ given by lateral juxtaposition - that is $\phi(a,b)=\phi_{0}(a)\phi_{m}(b)$.
@@
Proof. (I) Note from (4) that each group ${\mathsf{MD}}(n,n)$ is generated by the corresponding $s_{i}$’s and $r_{i}$’s. Their images at fixed $n$ are $1\otimes 1\otimes\cdots\otimes s\otimes 1\otimes\cdots\otimes 1$ ($s$ in the $i$-th position) and similarly for $r_{i}$.
@@
(4.6) Proposition. Wangian $\mathsf{MD}$} representations yield ${\mathsf{MD}}_{n}$ representations for all $n$ that are always semisimple over $\mathbb{C}$}. ∎
@@ 4.1. Notions of equivalence of representations/strict monoidal functors
This is precisely a monoidal natural equivalence between the functors $F,F^{\prime}$, in the sense of [MRT25].
@@
(II) Note that 2-equivalence does not imply $\infty$-equivalence [MRT25].
@@
(4.11) There are important (for efficient classification) weaker notions of equivalence (which in general do not imply even 2-equivalence), which might better be termed symmetries rather than equivalences. The point is that if there is a simple and manifest transformation that turns a solution into another solution then we need only classify one per orbit of this transformation. For more detail than we give here, see for example [MRT25, §2.2 et seq]. Otherwise, read on.
(4.12) As an example of such a symmetry we may extend Doiku-Smoktunowitz (DS) equivalence to $\mathsf{MD}$} in the obvious way: let $A\in\mathsf{Mat}^{N}(1,1)$ be such that $A\otimes A$ commutes with $R=F(r)$ and $S=F(s)$. Then $R^{\prime}=(A\otimes I_{N})R(A^{-1}\otimes I_{N})$ and $S^{\prime}=(A\otimes I_{N})S(A^{-1}\otimes I_{N})$ defines another functor $F^{\prime}$. In fact, the proof given in [MRT25, DS] for a single $R$-matrix goes through mutatis mutandis. It follows that the corresponding representations of ${\mathsf{MD}}_{n}$ are equivalent, i.e., $\infty$-equivalent.
@@ 4.2. On non-local but unentangled equivalences
The first of these is a solution despite being non-locally conjugated; and despite also that (as the second result shows) $R_{f}(p)$ does not commute with $\mathsf{h}\otimes\mathsf{h}$ in general (specifically unless $p^{2}=1$), so it is not of DS type. Note for comparison that $\mathsf{v}.R_{a}(p).\mathsf{v}$ is not a braid rep - more as one might expect.   This appears to raise several nice questions, but we will demote these to a later work.
@@
Since restriction to $R$ only, i.e. forgetting $S$, yields an ‘involutive’ braid representation - i.e. with $R^{2}=1$, we can start the classification by recalling the classification of such braid reps. This comes from Hietarinta’s rank-2 braid rep classification [Hietarinta1992] by restricting to the involutive cases.
Remark: It is interesting to note that of course this is the same as classifying ‘symmetric representations’ - functors $F:\Sigma\rightarrow\mathsf{Mat}^{2}$. In particular unitary ones have been classified up to $\infty$-equivalence in [LechnerPennigWood], but as we will see this is not the same as the classification up to $H_{1}$-equivalence that is needed here (see in particular Comment (4) below).
@@
We observe that two of these are obtained from charge-conserving (CC) solutions [MR1X] by adding ’glue’ i.e., some entries in the upper right. Two others are already charge-conserving. Thus convenient names for these are, respectively:
@@
(3) One can compare with the (elementary, partial) classification by $\Sigma_{2}$ character (or 2-equivalence class in the sense of 4.1 or [MRT25]): of course the trace of $R$ here is one of 4,2,0,-2,-4. Observe that the trivial $H_{1}$-class contains the 4 and -4 character classes. The two f-types and anti-slash contain the 2 and -2 character classes. And the two a-types together give the 0 character class.
(4) One can compare with the classification of unitary representations up to $\infty$-equivalence from [LechnerPennigWood]. These are indexed by pairs of Young diagrams such that the total degree is 2. Thus we have $(2,0)$, $(1^{2},0)$, $(1,1),(0,1^{2}),(0,2)$, corresponding to the trivial, f/a-slash type with $q=1$, the anti-slash type and overall sign changes.
Proof. One approach is simply to list all rank-2 braid reps up to the given equivalences and then impose the involutive condition. The ten ‘non-trivial’ varieties of such reps can be deduced from [Hietarinta1992] as follows.
@@ 5.1. Structure of representations: some preliminaries
The first two situations occur when the image of ${\mathbb{Z}}^{\binom{n}{2}}$ consists of diagonalisable matrices, and we can employ Theorems A.1 and A.4. Since the images of the ${\mathbb{Z}}^{\binom{n}{2}}$ generators are all similar to the Kronecker product of the matrix $X:=RS$ with identity matrix, we have: case (a) if $X$ has finite order and (b) if $X$ is diagonalisable but infinite order and (c) otherwise.
@@ 5.2. The MD representation in the f-glue case
(5.4) By the HoG Theorem B.2 we thus have the irreducible factors in all cases. Indeed the Bratelli diagram of the tower of semisimple quotients (a convenient summary of the irreducibles and their dimensions) is the usual (left-)truncated Pascal triangle.
@@ 5.3. The MD representation in the a-glue case
(5.7) So, back to the Wangian case: For general $n$ this is, up to isomorphism as above, the representation we get from Sutherland’s $gl_{q}(1|1)$ spin chain [Sutherland75] in the classical case (in the terminology of [MartinRittenberg92]). That is, it is a sum of all the two-row (+)(+-)-signed Young modules. Thus the structure is a direct sum of two copies of each of the $\Sigma_{n}$ irreps with a hook Young diagram label (see e.g. [MartinRittenberg92, DeguchiMartin92]). For example
@@
Proof. This follows immediately from the higher-on-glue theorem B.2. ∎
@@ 5.4. The MD representation in the antislash case
The eigenspace decomposition of $Y$ yields a decomposition of the representation of ${\mathsf{MD}}_{3}$ into $4$ irreducible representations of dimensions $1$ (twice) and $3$ (twice). The $1$-dimensional representations are trivial, and the $3$ dimensional representations are isomorphic copies of the irreducible representations found in Example A.10, with $X_{ij}$ eigenvalues $1,\lambda=-2z+1+2x,$ and $\lambda^{-1}=-2z+1-2x.$
@@
Set $W=\left[\begin{array}[]{cc}1&1\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 1&-1\end{array}\right]\otimes\left[\begin{array}[]{cc}1&-1\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 1&1\end{array}\right]$ and take $z=-\frac{(\lambda^{2}-2\lambda+1)}{4\lambda},x=\frac{(\lambda^{2}-1)}{4\lambda}.$ One easily checks that these satisfy $z^{2}-z+r^{2}=0$. Then we find that $WR^{\prime}W^{-1}=R$ and $WS^{\prime}W^{-1}=S$ so that the pairs $(R,S)$ and $R^{\prime},S^{\prime})$ are $2$-equivalent. Decomposing the ${\mathsf{MD}}_{3}$ representations yields $3$-equivalence.
On the other hand, one can show that corresponding functors are neither 1) locally equivalent or 2) DS-equivalent. Indeed, the only invertible $2\times 2$ matrices $A$ such that $A\otimes A$ commutes with both $R$ and $S$ are of the form: $A=\left[\begin{array}[]{cc}x&y\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr y&x\end{array}\right].$ But there are no such $A$ that map $(R,S)$ to $(R^{\prime},S^{\prime})$.

### v2

@@ 4. Mixed doubles representations
(4.2) The category $\mathsf{MD}$} is defined as follows. It is the strict monoidal category of mixed doubles groups ${\mathsf{MD}}_{n}$, analogous to the braid category or the Sym category. The object monoid is $(\mathbb{N}_{0},+)$, thus it is freely generated by the object 1. The monoidal product on morphisms is given by the inclusion $\phi:{\mathsf{MD}}_{m}\times{\mathsf{MD}}_{n}\rightarrow{\mathsf{MD}}_{m+n}$ given by lateral juxtaposition - that is $a\otimes b=\phi(a,b)=\phi_{0}(a)\phi_{m}(b)$ as defined in (2).
@@
Proof. (I) Note from (4) that each group ${\mathsf{MD}}(n,n)$ is generated by the corresponding $s_{i}$’s and $r_{i}$’s. Their images at fixed $n$ are $1\otimes\cdots\otimes 1\otimes s\otimes 1\otimes\cdots\otimes 1$ ($s$ in the $i$-th position) and similarly for $r_{i}$.
@@
(4.6) Proposition. Wangian $\mathsf{MD}$} representations yield ${\mathsf{MD}}_{n}$ representations for all $n$ that are always semisimple over the field $\mathbb{C}$}. ∎
@@ 4.1. Notions of equivalence of representations/strict monoidal functors
This is precisely a monoidal natural equivalence between the functors $F,F^{\prime}$, in the sense of [22].
@@
(II) Note that 2-equivalence does not imply $\infty$-equivalence [22].
@@
(4.11) There are important (for efficient classification) weaker notions of equivalence (which in general do not imply even 2-equivalence), which might better be termed symmetries rather than equivalences. The point is that if there is a simple and manifest transformation that turns a solution into another solution then we need only classify one per orbit of this transformation. For more detail than we give here, see for example [22, §2.2 et seq]. Otherwise, read on.
(4.12) As an example of such a symmetry we may extend Doiku-Smoktunowitz (DS) equivalence to $\mathsf{MD}$} in the obvious way: let $A\in\mathsf{Mat}^{N}(1,1)$ be such that $A\otimes A$ commutes with $R=F(r)$ and $S=F(s)$. Then $R^{\prime}=(A\otimes I_{N})R(A^{-1}\otimes I_{N})$ and $S^{\prime}=(A\otimes I_{N})S(A^{-1}\otimes I_{N})$ defines another functor $F^{\prime}$. In fact, the proof given in [22, 11] for a single $R$-matrix goes through mutatis mutandis. It follows that the corresponding representations of ${\mathsf{MD}}_{n}$ are equivalent, i.e., $\infty$-equivalent.
@@ 4.2. On non-local but unentangled equivalences
The first of these is a solution despite being non-locally conjugated; and despite also that (as the second result shows) $R_{f}(p)$ does not commute with $\mathsf{h}\otimes\mathsf{h}$ in general (specifically unless $p^{2}=1$), so it is not of DS type. Note for comparison that $\mathsf{v}.R_{a}(p).\mathsf{v}$ is not a braid representation - more as one might expect.   This appears to raise several nice questions, but we will demote these to a later work.
@@
Since restriction to $R$ only, i.e. forgetting $S$, yields an ‘involutive’ braid representation - i.e. with $R^{2}=1$, we can start the classification by recalling the classification of such braid reps. This comes from Hietarinta’s rank-2 braid rep classification [14] by restricting to the involutive cases.
Remark: It is interesting to note that of course this is the same as classifying ‘symmetric representations’ - functors $F:\Sigma\rightarrow\mathsf{Mat}^{2}$. In particular unitary ones have been classified up to $\infty$-equivalence in [17], but as we will see this is not the same as the classification up to $H_{1}$-equivalence that is needed here (see in particular Comment (4) below).
@@
We observe that two of these are obtained from charge-conserving (CC) solutions [20] by adding ’glue’ i.e., some entries in the upper right. Two others are already charge-conserving. Thus convenient names for these are, respectively:
@@
(3) One can compare with the (elementary, partial) classification by $\Sigma_{2}$ character (or 2-equivalence class in the sense of 4.1 or [22]): of course the trace of $R$ here is one of 4,2,0,-2,-4. Observe that the trivial $H_{1}$-class contains the 4 and -4 character classes. The two f-types and anti-slash contain the 2 and -2 character classes. And the two a-types together give the 0 character class.
(4) One can compare with the classification of unitary representations up to $\infty$-equivalence from [17]. These are indexed by pairs of Young diagrams such that the total degree is 2. Thus we have $(2,0)$, $(1^{2},0)$, $(1,1),(0,1^{2}),(0,2)$, corresponding to the trivial, f/a-slash type with $q=1$, the anti-slash type and overall sign changes.
Proof. One approach is simply to list all rank-2 braid reps up to the given equivalences and then impose the involutive condition. The ten ‘non-trivial’ varieties of such reps can be deduced from [14] as follows.
@@ 5.1. Structure of representations: some preliminaries
The first two situations occur when the image of ${\mathbb{Z}}^{\binom{n}{2}}$ consists of diagonalisable matrices. Since the images of the ${\mathbb{Z}}^{\binom{n}{2}}$ generators are all similar to the Kronecker product of the matrix $X:=RS$ with identity matrix, we have: case (a) if $X$ has finite order and (b) if $X$ is diagonalisable but infinite order and (c) otherwise.
@@ 5.2. The MD representation in the f-glue case
(5.4) By the HoG Theorem [2] we thus have the irreducible factors in all cases. Indeed the Bratelli diagram of the tower of semisimple quotients (a convenient summary of the irreducibles and their dimensions) is the usual (left-)truncated Pascal triangle.
@@ 5.3. The MD representation in the a-glue case
(5.7) So, back to the Wangian case: For general $n$ this is, up to isomorphism as above, the representation we get from Sutherland’s $gl_{q}(1|1)$ spin chain [29] in the classical case (in the terminology of [23]). That is, it is a sum of all the two-row (+)(+-)-signed Young modules. Thus the structure is a direct sum of two copies of each of the $\Sigma_{n}$ irreps with a hook Young diagram label (see e.g. [23, 10]). For example
@@
Proof. This follows immediately from the higher-on-glue theorem [2]. ∎
@@ 5.4. The MD representation in the antislash case
The eigenspace decomposition of $Y$ yields a decomposition of the representation of ${\mathsf{MD}}_{3}$ into $4$ irreducible representations of dimensions $1$ (twice) and $3$ (twice). The $1$-dimensional representations are trivial, and the $3$ dimensional representations are isomorphic copies of the same irreducible representation with $X_{ij}$ eigenvalues $1,\lambda=-2z+1+2x,$ and $\lambda^{-1}=-2z+1-2x.$
@@
Indeed, we find that these two representations are equivalent, with $\lambda$ as above. It is clear that they are not locally equivalent: any matrix of the form $A\otimes A$ commutes with $R^{\prime}$, so no local gauge choice can move $R^{\prime}$. Similarly, they are not DS-equivalent: there are no invertible matrices $T$ such that $I\otimes T$ conjugates the pair $(R^{\prime},S^{\prime})$ to $(R,S)$. On the other hand a combination of DS and local equivalences does yield an equivalence. We find that $T\otimes T$ where $T=\left[\begin{array}[]{cc}0&1\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 1&0\end{array}\right]$ commutes with both $R$ and $S$. Now take $A=\left[\begin{array}[]{cc}1&1\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr-1&1\end{array}\right]$ and define $M=(A\otimes A)(I\otimes T)=\left[\begin{array}[]{cccc}1&1&1&1\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 1&-1&1&-1\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr-1&-1&1&1\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr-1&1&1&-1\end{array}\right].$ Then $MRM^{-1}=R^{\prime}$ and $MSM^{-1}=S^{\prime}$, which is a composition of $\infty$-equivalences, the result follows.
(5.15) A similar approach show that case (b) with $\epsilon=-1$ $S=\left[\begin{array}[]{cccc}r&-y-1&y&-r\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr y&-r&r&-y-1\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr-y-1&r&-r&y\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr-r&y&-y-1&r\end{array}\right]$ is $\infty$-equivalent to case (5)
$R^{\prime},S^{\prime}]=\left[\left[\begin{array}[]{cccc}1&0&0&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 0&0&p&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 0&1/p&0&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 0&0&0&1\end{array}\right],\left[\begin{array}[]{cccc}1&0&0&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 0&0&\left(-2\,y-1-2\,r\right)p&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 0&{\frac{1}{\left(-2\,y-1-2\,r\right)p}}&0&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr 0&0&0&-1\end{array}\right]\right].$
In this case the same $A$ as above works, with $T=\left[\begin{array}[]{cc}p-1&p+1\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr p+1&p-1\end{array}\right].$